# Copositive criteria for a two-component dark matter model

### v1

@@ A.4 Cross section of $S_{i}S_{i}\to h_{i}h_{i}$ and $S_{i}S_{i}\to h_{1}h_{2}(i=1,2)$
We give the cross section of $S_{i}S_{i}\to h_{i}h_{i}$ and $S_{i}S_{i}\to h_{1}h_{2}$ $(i=1,2)$ in the part. We calculate the cross section with Calchep Belyaev:2012qa , and the expressions in the limit of $\sin\theta\to 0$ are given as follows.
@@
$+$ $\frac{8\lambda_{14}v_{0}^{2}(2\lambda_{14}v_{0}^{2}(m_{h_{1}}^{2}-s)-4m_{h_{1}}^{4}+s^{2})}{2m_{h_{1}}^{4}-3m_{h_{1}}^{2}s+s^{2}}\log(\frac{2m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{h_{1}}^{2})(s-4m_{1}^{2})}{s^{2}}}-1)}{2m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{1}^{2})(s-4m_{h_{1}}^{2})}{s^{2}}}+1)}))$ (45)
@@
$+$ $\frac{8\lambda_{13}v_{1}^{2}(2\lambda_{13}v_{1}^{2}(m_{h_{2}}^{2}-s)-4m_{h_{2}}^{4}+s^{2})}{2m_{h_{2}}^{4}-3m_{h_{2}}^{2}s+s^{2}}\log(\frac{2m_{h_{2}}^{2}+s(\sqrt{\frac{(s-4m_{h_{2}}^{2})(s-4m_{1}^{2})}{s^{2}}}-1)}{2m_{h_{2}}^{2}-s(\sqrt{\frac{(s-4m_{1}^{2})(s-4m_{h_{2}}^{2})}{s^{2}}}+1)}))$ (46)
@@
$+$ $\frac{8\lambda_{24}v_{0}^{2}(2\lambda_{24}v_{0}^{2}(m_{h_{1}}^{2}-s)-4m_{h_{1}}^{4}+s^{2})}{2m_{h_{1}}^{4}-3m_{h_{1}}^{2}s+s^{2}}\log(\frac{2m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{h_{1}}^{2})(s-4m_{2}^{2})}{s^{2}}}-1)}{2m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{2}^{2})(s-4m_{h_{1}}^{2})}{s^{2}}}+1)}))$ (47)
@@
$+$ $\frac{8\lambda_{23}v_{1}^{2}(2\lambda_{23}v_{1}^{2}(m_{h_{2}}^{2}-s)-4m_{h_{2}}^{4}+s^{2})}{2m_{h_{2}}^{4}-3m_{h_{2}}^{2}s+s^{2}}\log(\frac{2m_{h_{2}}^{2}+s(\sqrt{\frac{(s-4m_{h_{2}}^{2})(s-4m_{2}^{2})}{s^{2}}}-1)}{2m_{h_{2}}^{2}-s(\sqrt{\frac{(s-4m_{2}^{2})(s-4m_{h_{2}}^{2})}{s^{2}}}+1)}))$ (48)
@@
$+$ $\frac{s\sqrt{\frac{(s-4m_{1}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s}}}{2(-2m_{1}^{2}s(m_{h_{1}}^{2}+m_{h_{2}}^{2})+m_{1}^{2}(m_{h_{1}}^{2}-m_{h_{2}}^{2})^{2}+m_{h_{1}}^{2}m_{h_{2}}^{2}s+m_{1}^{2}s^{2})}$ (49)
@@
$+$ $\frac{s\sqrt{\frac{(s-4m_{2}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s}}}{2(-2m_{2}^{2}s(m_{h_{1}}^{2}+m_{h_{2}}^{2})+m_{2}^{2}(m_{h_{1}}^{2}-m_{h_{2}}^{2})^{2}+m_{h_{1}}^{2}m_{h_{2}}^{2}s+m_{2}^{2}s^{2})}$ (50)
According to the expressions of the cross section, in the limit of $\sin\theta\to 0$, the signs of $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$ can make difference on the terms including the odd orders of these couplings so that contribute to the cross section of $S_{i}S_{i}\to h_{i}h_{i}(i=1,2)$. Note that for the non-zero $\sin\theta$, the sign of $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$ can make also difference in the cross section by the combined contributions of these couplings such as $\lambda_{13}\lambda_{24}$.

### v4

@@ A.4 Cross section of $S_{i}S_{i}\to h_{i}h_{i}$ and $S_{i}S_{i}\to h_{1}h_{2}(i=1,2)$
We give the cross section of $S_{i}S_{i}\to h_{i}h_{i}$ and $S_{i}S_{i}\to h_{1}h_{2}$ $(i=1,2)$ in the part. We calculate the cross section with Calchep Belyaev et al. (2013), and the expressions in the limit of $\sin\theta\to 0$ are given as follows.
@@
$+$ $\frac{8\lambda_{14}v_{0}^{2}(2\lambda_{14}v_{0}^{2}(m_{h_{1}}^{2}-s)-4m_{h_{1}}^{4}+s^{2})}{2m_{h_{1}}^{4}-3m_{h_{1}}^{2}s+s^{2}}\log(\frac{2m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{h_{1}}^{2})(s-4m_{1}^{2})}{s^{2}}}-1)}{2m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{1}^{2})(s-4m_{h_{1}}^{2})}{s^{2}}}+1)})),$ (35)
@@
$+$ $\frac{8\lambda_{13}v_{1}^{2}(2\lambda_{13}v_{1}^{2}(m_{h_{2}}^{2}-s)-4m_{h_{2}}^{4}+s^{2})}{2m_{h_{2}}^{4}-3m_{h_{2}}^{2}s+s^{2}}\log(\frac{2m_{h_{2}}^{2}+s(\sqrt{\frac{(s-4m_{h_{2}}^{2})(s-4m_{1}^{2})}{s^{2}}}-1)}{2m_{h_{2}}^{2}-s(\sqrt{\frac{(s-4m_{1}^{2})(s-4m_{h_{2}}^{2})}{s^{2}}}+1)})),$ (36)
@@
$+$ $\frac{8\lambda_{24}v_{0}^{2}(2\lambda_{24}v_{0}^{2}(m_{h_{1}}^{2}-s)-4m_{h_{1}}^{4}+s^{2})}{2m_{h_{1}}^{4}-3m_{h_{1}}^{2}s+s^{2}}\log(\frac{2m_{h_{1}}^{2}+s(\sqrt{\frac{(s-4m_{h_{1}}^{2})(s-4m_{2}^{2})}{s^{2}}}-1)}{2m_{h_{1}}^{2}-s(\sqrt{\frac{(s-4m_{2}^{2})(s-4m_{h_{1}}^{2})}{s^{2}}}+1)})),$ (37)
@@
$+$ $\frac{8\lambda_{23}v_{1}^{2}(2\lambda_{23}v_{1}^{2}(m_{h_{2}}^{2}-s)-4m_{h_{2}}^{4}+s^{2})}{2m_{h_{2}}^{4}-3m_{h_{2}}^{2}s+s^{2}}\log(\frac{2m_{h_{2}}^{2}+s(\sqrt{\frac{(s-4m_{h_{2}}^{2})(s-4m_{2}^{2})}{s^{2}}}-1)}{2m_{h_{2}}^{2}-s(\sqrt{\frac{(s-4m_{2}^{2})(s-4m_{h_{2}}^{2})}{s^{2}}}+1)})),$ (38)
@@
$+$ $\frac{s\sqrt{\frac{(s-4m_{1}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s}}}{2(-2m_{1}^{2}s(m_{h_{1}}^{2}+m_{h_{2}}^{2})+m_{1}^{2}(m_{h_{1}}^{2}-m_{h_{2}}^{2})^{2}+m_{h_{1}}^{2}m_{h_{2}}^{2}s+m_{1}^{2}s^{2})},$ (39)
@@
$+$ $\frac{s\sqrt{\frac{(s-4m_{2}^{2})(m_{h_{1}}^{4}-2m_{h_{1}}^{2}(m_{h_{2}}^{2}+s)+(m_{h_{2}}^{2}-s)^{2})}{s}}}{2(-2m_{2}^{2}s(m_{h_{1}}^{2}+m_{h_{2}}^{2})+m_{2}^{2}(m_{h_{1}}^{2}-m_{h_{2}}^{2})^{2}+m_{h_{1}}^{2}m_{h_{2}}^{2}s+m_{2}^{2}s^{2})}.$ (40)
According to the expressions of the cross-section, in the limit of $\sin\theta\to 0$, the signs of $\lambda_{13}$$\lambda_{23}$$\lambda_{14}$ and $\lambda_{24}$ can make difference on the terms including the odd orders of these couplings so that contribute to the cross-section of $S_{i}S_{i}\to h_{i}h_{i}(i=1,2)$. Note that for the non-zero $\sin\theta$, the sign of $\lambda_{13}$$\lambda_{23}$$\lambda_{14}$ and $\lambda_{24}$ can make also difference in the cross section by the combined contributions of these couplings such as $\lambda_{13}\lambda_{24}$.
Figure 7: Results of viable parameter space satisfying the relic density and the direct detection constraint for case (1). The picture (a) shows the allowed values for dark matter mass $S_{1}$(red points) and $S_{2}$ (blue points), and the black line represents the upper bound according to the latest LZ results. The picture (b) shows the viable parameter space for $m_{1}-\lambda_{14}$, where bubbles with different colors correspond to the fraction of $S_{1}$ defined by $\Omega_{1}/(\Omega_{1}+\Omega_{2})$. The picture (c) shows the viable parameter space for $m_{2}-\lambda_{24}$, where bubbles with different colors correspond to the fraction of $S_{2}$ defined by $\Omega_{2}/(\Omega_{1}+\Omega_{2})$. The picture (d) shows the allowed value for $\lambda_{13}-\lambda_{23}$, where bubbles with different colors represent the viable value for $\lambda_{12}$.
Figure 8: Results of viable parameter space satisfying the relic density and the direct detection constraint for case (2). In (b),(c),(d), the colored points satisfying copositive criterial while the colored bubbles are not, which are labeled by ”copo” and ”no copo”.
Figure 9: Results of viable parameter space satisfying the relic density and the direct detection constraint for case (3). In (b),(c),(d), the colored points satisfying copositive criterial while the colored bubbles are not, which are labeled by ”copo” and ”no copo”.
Figure 10: Results of viable parameter space satisfying the relic density and the direct detection constraint for case (4). In (b),(c),(d), the colored points satisfying copositive criterial while the colored bubbles are not, which are labeled by ”copo” and ”no copo”.
A.5 Random scan results
In this part, we perfrom a random scan on the chosen parameter space with:
$m_{1,2}\subset[1\mathrm{TeV},4\mathrm{TeV}],|\lambda_{12}|\subset[10^{-4},3.14],|\lambda_{13}|\subset[10^{-4},3.14],$
$|\lambda_{14}|\subset[10^{-4},3.14],|\lambda_{23}|\subset[10^{-4},3.14],|\lambda_{24}|\subset[10^{-4},3.14].$
while we fix $\sin\theta=0.01,m_{h_{2}}=1.5$ TeV and $v_{1}=2$ TeV for simplicity. We consider the copositive criteria, dark matter relic density and direct detection constraint on the model for the four different cases, and the reuslts are given in Fig. 7 to Fig. 10.
For the case (1), the copositive criterial makes little difference in the parameter space, and we show the viable parameter space satisfying the relic density and the direct detection constraint in Fig. 7. Both $m_{1}$ and $m_{2}$ can take values ranging from [1 TeV, 4 TeV] as we can see in Fig. 7(a). According to (b) and (c), a larger $\lambda_{14}$ ($\lambda_{24}$) can always contribute to a stronger interaction rate so that the fraction of $S_{1}$ ($S_{2}$) is smaller. The direct detection constraint limits a small $\lambda_{14}$ and $\lambda_{24}$, and a large $\lambda_{13}$ as well as $\lambda_{23}$ are hence demanded under relic density constraint. As we can see in Fig. 7(d), most of the points lie in the upper right region for $(\lambda_{13},\lambda_{23})$. In addition, $\lambda_{12}$ is related to the conversion processes between $S_{1}$ and $S_{2}$, which is less constrained.
For the case (2), case (3) and case (4), we consider the copositive criterial on the parameter space, and we show the viable parameter space satisfying the relic density and the direct detection constraint in Fig. 8, Fig. 9 and Fig. 10, where the colored points satisfying copositive criterial while the colored bubbles are not, which are labeled by ”copo” and ”no copo”. We have similar conclusion for the case (1), but when considering copositive criterial, the parameter space is more constrained and only a few points ”survived” according to Fig. 8 to Fig. 10. In addition, for the different copositive criterials, the allowed parameter space are different.